# Helson’s conjecture for smooth numbers

### v2

@@ Abstract.
Let $\Psi(x,y)$ denote the count of $y$-smooth numbers below $x$ and $P(n)$ denote the largest prime factor of $n$. We prove that for $f$ a Steinhaus random multiplicative function, the partial sums over $y$-smooth numbers enjoy better than squareroot cancellation, in the sense that
@@
uniformly for $(\log x)^{30}\leqslant y\leqslant x$. Our bounds are quantitative and give a large saving when $y$ isn’t too close to $x$.
@@ 1.1. Motivation and statement of results
The analogous character and zeta upper bounds are also proved by Harper in [harpertypical] using a remarkable derandomisation method. More precisely, Harper explicitly relates these moments of character sums and zeta sums to (a certain stage in the proof of) the random multiplicative case. As is to be expected, the length of the sum in that case depends on the parameters $r$ and $T$ that determine the size of the family of characters. For example, it is shown that for any large prime $r$ and for all $x\leqslant r$ with $x,\frac{r}{x}\to+\infty$, one has
@@
The key idea behind Harper’s theorem is to notice that the partial sums are connected to a phenomenon known as critical Gaussian multiplicative chaos (GMC). To briefly elaborate, we define $F_{x}(s)\coloneqq\prod_{p\leqslant x}\bigl(1-\frac{f(p)}{p^{s}}\bigr)^{-1}$ to be the truncated Euler product. By a non-trivial conditioning argument and an application of Plancherel’s identity, the proof works by first showing that
@@
In this paper, our attention lies on what happens when $\mathcal{A}$ is the set of $y$-smooth numbers, i.e., all elements in $\mathcal{A}$ only have prime factors $p\leqslant y$. Here and throughout the paper, let $P(n)$ denote the largest prime factor of $n$, so that $\Psi(x,y)=\#\{n\leqslant x:P(n)\leqslant y\}$ denotes the count of $y$-smooth numbers up to $x$. Note that the case where $y=x$ corresponds to the full sum, for which we know from Harper[HarperLow] that we have better than squareroot cancellation. Analogously to the previous cases, one might wonder if there is a certain transition range where the “better than squareroot” cancellation disappears. Our main theorem suggests that this may not be the case.
@@
Suppose that $(\log x)^{30}\leqslant y\leqslant x$. Then we have
uniformly for $y$ in this range. Quantitative bounds can be found in Theorems 1.3 and LABEL:t:small_u_upper_bound.
@@
$E(x,x^{\varepsilon},h)\ll\frac{1}{x}\int_{\begin{subarray}{c}|t|\leqslant T\\ t\not\approx 0\end{subarray}}\Bigl|\sum_{\begin{subarray}{c}n\leqslant x+h\\ P(n)\leqslant x^{\varepsilon}\end{subarray}}n^{-it}\Bigr|\,dt.$
@@
In particular, since $\Psi(x,x^{\varepsilon})\asymp_{\varepsilon}x$, combining Theorem 1.1 and (1.2) (with $A=2$) gives the bound $E(x,x^{\varepsilon},\sqrt{x})=o\big(\sqrt{\Psi(x,x^{\varepsilon})/x}\big)=o(1)$, which (very roughly) gives an asymptotic for the count in intervals of length $h=\sqrt{x}$. In fact, our quantitative result for $y$ in this range, Theorem LABEL:t:small_u_upper_bound, would suggest further that this strategy could show the existence of $x^{\varepsilon}$-smooth numbers in intervals of length $h\gg\sqrt{x}/(\log\log x)^{1/4}$, similarly to how the analogous conjecture in [harpertypical] gives Möbius cancellation in sums of length $h\gg\sqrt{x}/(\log\log x)^{1/4}$.
Based on Theorem 1.1, it seems reasonable to conjecture that “better than squareroot cancellation” should hold uniformly for all $y\leqslant x$.
Conjecture 1.2.
Theorem 1.1 holds uniformly for any $2\leqslant y\leqslant x$.
It would be nice if one could solve Conjecture 1.2 completely, though our main theorem covers what we believe to be the most interesting range. Analogously to (1.1), we always work by connecting the partial sums to the Euler product. To illustrate how the analysis changes for smaller $y$, in the most extreme case, when $y=2$, we have
$\mathbb{E}\Bigl|\sum_{\begin{subarray}{c}n\leqslant x\\ P(n)\leqslant 2\end{subarray}}f(n)\Bigr|=\mathbb{E}\Bigl|\sum_{n\leqslant\lfloor\frac{\log x}{\log 2}\rfloor}f(2)^{n}\Bigr|=\int_{0}^{1}\Bigl|\sum_{n\leqslant\lfloor\frac{\log x}{\log x}\rfloor}\mathrm{e}(n\theta)\Bigr|\,d\theta\ll\log\log x,$
which is very small relative to $\Psi(x,2)^{1/2}\asymp(\log x)^{1/2}$. This provides evidence towards our conjecture, and also points out that for small $y$, the problem is related to finding cancellation in exponential sums over lattices. The exponential sum estimate becomes challenging once $y\gg(\log x)^{1+o(1)}$, and new ideas are needed to cover this range.
@@
In this paper, we observe two distinct sources for the “better than squareroot cancellation” phenomena in Theorem 1.1; the size of $y$ being key to this analysis. First of all, for $y$ being close to $x$, one should expect that the critical GMC phenomenon uncovered by Harper [HarperLow] should give better than squareroot cancellation. We find this to be the case, and moreover, we are able to use critical GMC phenomenon to prove better than squareroot cancellation even when $y$ is as small as $e^{(\log\log x)^{2}}$, say. We outline how this is done in Section LABEL:s:_large_y_explained, and a precise statement is given in Theorem LABEL:t:small_u_upper_bound. The critical GMC saving in our case specifically comes from exploiting the randomness of $f(p)$ on the small primes (to be precise, primes smaller than roughly $e^{\frac{1}{1-\alpha}}$ for $\alpha=\alpha(x,y)$ the saddle point, see Section LABEL:s:smooths_intro for details), and, as one may expect, the saving is always quite small (like a power of $\log\log x$, or smaller).
@@
Given the previous works [HarperLow, Xu, Caichshort], one may not expect to observe cancellation that is larger than some power of $\log\log x$ for any reasonably large $y$. However, we find that when $y$ is smaller than roughly $x^{\frac{1}{\log\log x}}$, we can obtain an unexpectedly large saving (actually coming from the primes closer to $y$). In short, this saving comes from the fact that the dominant contribution to certain Euler product expectations comes from highly unlikely events (we explain this in detail in Section LABEL:s:outline_of_moderate_y). The flavour of this might remind some readers of the supercritical GMC case, although this is not exactly what we observe, since our Euler products are shifted off the half-line and are no longer approximated by the exponential of a log-correlated Gaussian field. This novel phenomenon allows us to prove the following:
Theorem 1.3.
Let $u=\frac{\log x}{\log y}$ and fix $\varepsilon>0$. For $e^{(\log\log x)^{1+\varepsilon}}\leqslant y\leqslant x^{\frac{1}{(\log\log x)^{1+\varepsilon}}}$, we have

### v3

@@ Abstract.
Let $\Psi(x,y)$ denote the count of $y$-smooth numbers below $x$ and $P(n)$ denote the largest prime factor of $n$. We prove that for $f$ a Steinhaus random multiplicative function, the partial sums over $y$-smooth numbers always enjoy better than squareroot cancellation, in the sense that
@@
uniformly on the entire range $2\leqslant y\leqslant x$. The bounds are quantitative and give a large saving when $y$ isn’t too close to $x$.
@@ 1.1. Motivation and statement of results
The analogous character and zeta upper bounds are also proved by Harper in [harpertypical] using a remarkable comparison method. More precisely, Harper explicitly relates these moments of character sums and zeta sums to (a certain stage in the proof of) the random multiplicative case. As is to be expected, the length of the sum in that case depends on the parameters $r$ and $T$ that determine the size of the family of characters. For example, it is shown that for any large prime $r$ and for all $x\leqslant r$ with $x,\frac{r}{x}\to+\infty$, one has
@@
The key idea behind Harper’s theorem is to notice that the partial sums are connected to a phenomenon known as critical Gaussian multiplicative chaos (GMC). To briefly elaborate, we define $F_{x}(s)\coloneqq\prod_{p\leqslant x}\bigl(1-\frac{f(p)}{p^{s}}\bigr)^{-1}$ to be the truncated Euler product. By a non-trivial conditioning argument and an application of Parseval’s identity, the proof works by first showing that
@@
In this paper, our attention lies on what happens when $\mathcal{A}$ is the set of $y$-smooth numbers, i.e., all elements in $\mathcal{A}$ only have prime factors $p\leqslant y$. Here and throughout the paper, let $P(n)$ denote the largest prime factor of $n$, so that $\Psi(x,y)=\#\{n\leqslant x:P(n)\leqslant y\}$ denotes the count of $y$-smooth numbers up to $x$. Note that the case where $y=x$ corresponds to the full sum, for which we know from Harper [HarperLow] that we have better than squareroot cancellation. Analogously to the previous cases, one might wonder if there is a certain transition range where the “better than squareroot” cancellation disappears. Our main theorem shows that this is not the case:
@@
Uniformly for $2\leqslant y\leqslant x$, we have
$\mathbb{E}\Big|\sum_{\begin{subarray}{c}1\leqslant n\leqslant x\\ P(n)\leqslant y\end{subarray}}f(n)\Big|=o\left(\Psi(x,y)^{1/2}\right).$
Quantitative bounds can be found in Theorem 1.2 (for moderately sized $y$), Theorem LABEL:t:small_u_upper_bound (for $y$ close to $x$), and Theorem LABEL:t:_small_y (for small $y$).
@@
$E(x,x^{\varepsilon},h)\ll\frac{1}{x}\int_{\begin{subarray}{c}|t|\leqslant x/h\\ t\not\approx 0\end{subarray}}\Bigl|\sum_{\begin{subarray}{c}n\leqslant x+h\\ P(n)\leqslant x^{\varepsilon}\end{subarray}}n^{-it}\Bigr|\,dt.$
@@
In particular, since $\Psi(x,x^{\varepsilon})\asymp_{\varepsilon}x$, combining Theorem 1.1 and (1.2) gives the bound $E(x,x^{\varepsilon},\sqrt{x})=o\big(\sqrt{\Psi(x,x^{\varepsilon})/x}\big)=o(1)$, which (very roughly) gives an asymptotic for the count in intervals of length $h=\sqrt{x}$. In fact, our quantitative result for $y$ in this range, Theorem LABEL:t:small_u_upper_bound, would suggest further that this strategy could show the existence of $x^{\varepsilon}$-smooth numbers in intervals of length $h\gg\sqrt{x}/(\log\log x)^{1/4}$, similarly to how the analogous conjecture in [harpertypical] gives Möbius cancellation in sums of length $h\gg\sqrt{x}/(\log\log x)^{1/4}$.
@@
In this paper, we observe three distinct sources for the “better than squareroot cancellation” phenomena in Theorem 1.1; the size of $y$ being key to this analysis. First of all, for $y$ close to $x$, one should expect that the critical GMC phenomenon uncovered by Harper [HarperLow] should give better than squareroot cancellation. We find this to be the case, and moreover, we are able to use critical GMC phenomenon to prove better than squareroot cancellation even when $y$ is as small as $e^{(\log\log x)^{5/3+\varepsilon}}$. We outline how this is done in Section LABEL:s:_large_y_explained, and a precise statement is given in Theorem LABEL:t:small_u_upper_bound. The critical GMC saving in our case specifically comes from exploiting the randomness of $f(p)$ on the small primes (to be precise, primes smaller than roughly $e^{\frac{1}{1-\alpha}}$ for $\alpha=\alpha(x,y)$ the saddle point, see Section LABEL:s:smooths_intro for details), and, as one may expect, the saving is always quite small (like a power of $\log\log x$, or smaller).
@@
Given the previous works [Caichshort, HarperLow, Xu], one may not expect to observe cancellation that is larger than some power of $\log\log x$ for any reasonably large $y$. However, we find that when $y$ is smaller than roughly $x^{\frac{1}{\log\log x}}$, we can obtain an unexpectedly large saving (coming from the primes closer to $y$). In short, this saving comes from the fact that the dominant contribution to our Euler product expectations comes from highly unlikely events (we explain this in detail in Section LABEL:s:outline_of_moderate_y). The flavour of this might remind some readers of the supercritical GMC case, although this is not exactly what we observe, since our Euler products are shifted off the half-line and are no longer approximated by the exponential of a log-correlated Gaussian field. This allows us to prove the following:
Theorem 1.2 (Moderate $y$).
Let $u=\frac{\log x}{\log y}$ and fix $\varepsilon>0$. For $(\log x)^{1+\varepsilon}\leqslant y\leqslant x^{\frac{1}{(\log\log x)^{1+\varepsilon}}}$, we have